OpenStudy (anonymous):

John is 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages.

OpenStudy (anonymous):

J=3S (J+5)=2(S+5) solve for J and S

OpenStudy (anonymous):

3S+5=2S+10 Solve for S and then calculate J